OpenStudy (freckles):

\[x^3+3x^2+3x+1 \\ x^3+1+3x^2+3x \\ (x+1)(x^2-x+1)+3x(x+1) \\ (x+1)(x^2-x+1+3x) \\ (x+1)(x^2+2x+1) \\ (x+1)(x+1)^2 \\ (x+1)^3 \\ \text{ So } 343=1331_6 \\ \] Thought it was weird in base 6 this number had a factorization like this. I get 7 and 7^2... But 6. I wonder why 6 is unique here amongst the non-factors of 343. Or I don't know if it is unique or not. I have just did 343 in base whatever on wolfram all night last night looking for numbers that would give this kinda factorization.

OpenStudy (freckles):

My goal is to somehow find a way to write mnmnmnmnm.... (odd amount of digits and this symmetry all the way through) in some base number x such that the number can factored in a similar manner above.

OpenStudy (misty1212):

HI \(\color\magenta\heartsuit\)

OpenStudy (freckles):

Hi!

OpenStudy (rational):

@Kainui

OpenStudy (freckles):

By the way I do warn you guys I don't know if this can be done. For some reason I assume it can be.

OpenStudy (freckles):

\[34343=9:14:0_{61} \\ 9x^2+14x+0x \\ x(9x+14) \\ \text{ Like I don't like 61 much because it is a factor of 34343 }\] It would be cool to find another odd ball like we did above for 343.

OpenStudy (mathmate):

If I understand correctly the phenomenon, I wonder if it wasn't because 7=6+1 If we try 6=5+1, we have \(6^3=216_{10}=1331_5\)

OpenStudy (freckles):

hmm... that's neat

OpenStudy (freckles):

so \[5^3=125=1331_4 \]

OpenStudy (mathmate):

It also works for 9: \(9^3=729=1331_8\) :)

OpenStudy (freckles):

\[n^3=1331_{n-1} \text{ for integer n}\] I wonder if we can prove that by induction.

OpenStudy (freckles):

n would have to be bigger than ...

OpenStudy (mathmate):

Tag @ganeshie8 , he can do anything!

OpenStudy (freckles):

well there is a 3 in there so n would have to be bigger than or equal to 5

OpenStudy (mathmate):

Yep!

OpenStudy (mathmate):

Or by mod.

OpenStudy (rational):

@Marki

OpenStudy (anonymous):

hmm im like my brain crashed a bit hahaha :\ somehow

OpenStudy (anonymous):

i've seen this ... let me get my memory back

OpenStudy (freckles):

\[\text{ Assume } n^3=1331_{n-1} \\ \text{ for integer } n \ge 5 \\ \text{ We want to show } (n+1)^3=1331_{n}\] \[n^3+3n^2+3n+1=1331_{n-1}+?????\] Trying to figure out what to do with the 3n^2+3n+1 part

OpenStudy (freckles):

by the way while I think @rational were you assigned that 343 /34343 / so on thing?

OpenStudy (freckles):

Like was that an assignment ? And if it was there your teacher ever tell you the way they would have done it?

OpenStudy (freckles):

did* (not there)

OpenStudy (anonymous):

i think this is too smart

OpenStudy (freckles):

If it is too smart for you @marki then it is too smart for me :(

OpenStudy (freckles):

You are a genius.

OpenStudy (anonymous):

smart = think about = solvable = not out from space = u are smart to get such idea also dont get such idea's of hmmm :|

OpenStudy (rational):

\[(1331)_{n-1} = (n-1)^3 + 3(n-1)^2 + 3(n-1) + 1 = n^3\]

OpenStudy (freckles):

neat neat!

OpenStudy (rational):

@freckles that 34343.. was a very hard problem which i have been working on and off during my free time..

OpenStudy (rational):

@Marki and @Kainui also know about that problem more than me

OpenStudy (freckles):

Like even digit numbers like are way easy. http://openstudy.com/users/freckles#/updates/54a9a050e4b054f0c3ba5a87

OpenStudy (freckles):

I been plugging with this symmetry and this way of factoring for a bit. I only kinda got the different base idea from @ganeshie8.

OpenStudy (freckles):

http://openstudy.com/users/freckles#/updates/54ab17aae4b0b83d9dca4479 When he posted here about it,

OpenStudy (anonymous):

freckles we already solved for these cases:- if number of digits of the form 3n and 3n+1

OpenStudy (anonymous):

i cant remember where its been showing all work

OpenStudy (freckles):

for that induction problem?

OpenStudy (freckles):

I think @rational actually just proved it

OpenStudy (freckles):

and I feel like an idiot for not seeing it :p

OpenStudy (anonymous):

ugh no

OpenStudy (rational):

this changing base thingy is giving me some hope for that 34343.. xD @FibonacciChick666 is good with NT too

OpenStudy (freckles):

@zarkon and @sithgiggles is that how you spell his name (it isn't showing up for me-that sithgiggles guy)

OpenStudy (rational):

@SithsAndGiggles

OpenStudy (anonymous):

know what i'll work in this -.- i cant run away :P

OpenStudy (fibonaccichick666):

Would love to help, but I honestly have no clue here

OpenStudy (freckles):

so 6 I think won't help at all for the other numbers because none of the other numbers in the sequence mentioned by @rational can be written as some number cube

OpenStudy (freckles):

or you know the relationship between 6 and 343 will not help us at all*

OpenStudy (freckles):

ok well I have another way to proceed ... I wonder if writing a number like this is : \[mnmnmnmnmnm....=r:0_x\] is doable where r is a number between 0 and x (including 0 but not x) So we have a number in base 10 that we want to write in base x. And this number in base x only has two digits where the one digit is 0.

OpenStudy (freckles):

for example: \[34343_{10}=61:0_{563}\]

OpenStudy (fibonaccichick666):

Here is my guess, the prime factorization of 6 is 2*3 now, 4 is just 2*2 so the digits of 343434.... can be made using only the numbers in the prime factorization of 6

OpenStudy (freckles):

but that is cheating I think

OpenStudy (freckles):

not you @FibonacciChick666 I was talking about myself (and the cheating thing)

OpenStudy (fibonaccichick666):

lol it's cool

OpenStudy (freckles):

Like I think the (r:0)_x thing works only if you know a factor of the number given but it would cool to write it in that base number without even assuming it was a factor but I don't know if that is possible

OpenStudy (anonymous):

@freckles here is the main topic http://openstudy.com/study#/updates/53ecad4be4b01789aba50084

OpenStudy (rational):

that looks neat :) here are some easy observations about the sequence 34343... : 1) starting from third term, every other third term is divisible by 3 because the sum of digits is divisible by 3 2) starting from first term, every other third term is divisible by 7 for another cool reason

OpenStudy (anonymous):

yep yep we showed that ealier our problem with when number of digits is of the form 3n+2 like 34343 34343434343 34343434343434343 ..............

OpenStudy (anonymous):

but i want to forget this right now and work with bases like what u asked @freckles

OpenStudy (freckles):

I might have to take a break from this because at the moment I feel like I don't know where to go.

OpenStudy (freckles):

I really like this though. People like us working on something new I think. I think is the new.

OpenStudy (anonymous):

well i like myself working on this , hmm i dont think other ppl really care :P

OpenStudy (freckles):

Well I like care. I like that you are thinking about it with me.

OpenStudy (anonymous):

^_^

OpenStudy (rational):

\[\large 34343434343_{10}= 3713:0_{9249511}\] essentially you have rephrased the problem, thats all xD

OpenStudy (anonymous):

whats this xD

OpenStudy (rational):

\[\large 34343434343_{10}= 1:0_{34343434343}\]

OpenStudy (anonymous):

?

OpenStudy (rational):

\[\large 34343434343_{10}= 10_{34343434343}\]

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

oh no :P

OpenStudy (anonymous):

oh God xD i'll fell down

OpenStudy (rational):

\[\large (343)_{10} = (7\times 49)_{10} = (70)_{49}\]

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

solved omg i still feels like WOW

OpenStudy (rational):

still not clear ha ?

OpenStudy (anonymous):

NO ITS SO CLEAR !

OpenStudy (freckles):

wait but how are you figuring out the base number such that it can be written as a 2 digit number with 1's digit as 0

OpenStudy (freckles):

\[\large 34343434343_{10}= 10_{34343434343} \text{ this one is most interesting }\]

OpenStudy (freckles):

wait

OpenStudy (rational):

there is no way to find that afaik :P wolf uncle is helping me

OpenStudy (freckles):

\[34343=10_{34343}\]

OpenStudy (anonymous):

so since all of the sequence of the form 10 base 34343... that means (2*5) in base 3434343... means all composite :3

OpenStudy (freckles):

so wait wait if they all can be written like this...

OpenStudy (freckles):

oh no

OpenStudy (rational):

lol its not gona help us

OpenStudy (freckles):

doesn't help because that one digit is 1

OpenStudy (rational):

;)

OpenStudy (freckles):

so r:0_x where r is [0,1) U (1,x)

OpenStudy (rational):

there is no free lunch haha!

OpenStudy (anonymous):

hmm ok feeling dull a bit

OpenStudy (anonymous):

what does that means ?

OpenStudy (freckles):

I'm beginning to think this approach will not work

OpenStudy (anonymous):

brb 5 mnts

OpenStudy (freckles):

But who knows that could be me giving up early

OpenStudy (anonymous):

ok ok ok

OpenStudy (anonymous):

im quit bored :P